# Rui Eduardo Lopes, Student Member, IEEE, Susana Sargento, Senior Member, IEEE

### v1

@@ III-E2 In wireless context
  • Interference from other devices communicating on the same spectrum frequencies. This issue is more pronounced when unlicensed spectrum (e.g. 2.4 GHz) is used since its usage is uncontrolled and prone to additional noise from other uncontrollable entities.
  • Shadowing and attenuation from other physical obstacles and phenomena (e.g. reflection, absorption, diffraction) [110].
@@
  • Bounding an end-to-end latency or its jitter by integrating scheduling on transmission by each Access Point (AP).
  • Integration of reliability techniques while considering and taking maximum advantage of the main different channel characteristics.
@@
Achieving determinism within a wireless communication requires that all these challenges are properly addressed.
While fading, interference, shadowing and attenuation are mostly unavoidable without changes to the physical environment, interference can be minimised. The usage of other parts of the spectrum, possibly licensed, with wider channels, and frequencies in multiple links working in conjunction (multi link operation) can further decrease the overall noise within the physical layer and therefore decrease the jitter of latencies, and increase the determinism of the communications [103]. In addition, a periodic analysis of the spectrum usage can support a decision on the best frequencies to be used.
Regarding the unpredictable time to access the wireless medium, existing listen-before-talk solutions such as Carrier-Sense Multiple Access (CSMA) consider that in the event of a collision, before re-transmission a random amount of time shall be waited. This creates an unpredictable, non-deterministic jitter on the transmission of information. Other channel access solutions must be considered in order to achieve higher determinism on the transmission of the frames through both a wired or wireless medium, such as dividing the channel per frequency, time or both — Frequency-Division Multiple Access (FDMA), TDMA and  Orthogonal FDMA (OFDMA), respectively — or adding scheduling algorithms to listen-before-talk solutions [111]CSMA-Collision Avoidance (CSMA-CA). In alternative, solutions that consider collisions to be solved at the receptor side — e.g.  Non-orthogonal Multiple Access (NOMA) — can also be considered.
Another promising solution is a coordination and scheduling of the access to the wireless medium to ensure not fairness or throughput — the usual concerns — but the fulfillment of a specific time of transmission, ensuring more determinism than the previous approaches [103].
Within wireless mediums, the usage of the IEEE 802.11be (Wi-Fi 7) concept of restricted Target Wake Time (rTWT) — originally introduced for power saving — can further decrease conflicts during transmission. By defining when a certain station cannot transmit — and therefore any transmission ends before the start of this time — it is possible to further decrease collisions between critical and less critical frames being transmitted simultaneously.
@@
Regarding the bounding of end-to-end latency and jitter, the integration of shaping techniques, namely TAS/802.11Qbv or ATS is possible. Solutions for integration include mapping the scheduling policies to each transmission queue within an AP and by adding to each frame fields with the flow identification and characteristics (delay bound, maximum packet size) to map the scheduling policies to the 802.11be MAC. The overall scheduling requires a full knowledge of the network topology.
@@
Finally, regarding the resource management, the need for some sort of centralization in decision making for the configuration of the network nodes — APs and switches — is supported both by TSN and DetNet standards. TSN IEEE 802.1Qcc standard define that 2 entities — the CUC and CNC — are responsible for monitoring and defining resource allocation and configurations [114], while the DetNet Reference Architecture [92] considers the existence of a Controller Plane responsible for “topology discovery and maintenance, packet route selection and instantiation or path fail-over mechanisms” [115]. An architecture should therefore consider the existence at least of a controller that configures and monitors the network nodes and the information flow.
All the previous points can benefit from more flexibility and performance with a deployment using FPGAs [116]. FPGAs, by allowing the reconfiguration of the hardware to fit the needs of the application(s), can support the deployment of new wireless APs that combine the previous techniques accelerated by hardware. Such acceleration is essential to ensure the processing times are sufficiently small to allow determinism in the overall communication.
@@
  • Timer: used by the scheduler, is synchronized by an external Pulse Per Second (PPS) signal.
  • Scheduler: operating at level 2, decides which frames should be sent and when are they due.
  • Custom MAC: contains information for controlling frames, with changes from the original 802.11 access scheme to avoid collision between real-time and non-real-time frames and support the synchronization between clocks.
@@
A promising solution that however lacks full testing in mobility scenarios, integration with TSN and integration with other wireless standards such as 5G-NR. Such integration with cellular communication technologies is possible. 3GPP Releases 16-18 defends that, ensuring ultra reliable and low latency communications start at the radio level. TSN standards should be added to the overall architecture, with parts of the 5G core acting as TSN bridges [114], one per 5G User Plane Function (UPF[118]. This allows a transparent integration of 5G systems with a TSN system but require a set of translators between the 5G control plane and TSN standards [119].
The configuration of the TSN network is based on the 5G configuration, with the mapping between the TSN network configurations [120] — such as priorities or TAS gate opening times — and 5G QoS parameters — such as the 5G QoS Identifier (5QI), a scalar that represents characteristics such as [118, 121]:
  • Resource Type: Guaranteed Bit Rate (GBR), Delay Critical GBR (DC-GBR) or non-GBR, where the first two can pre-allocate resources.
  • Priority Level: indicates a priority in scheduling resources. Can be used to prioritize the allocation of resources to certain flows.
  • Delay Budget: an upper bound of the time a packet can spend inside a 5G system (between the User Equipment (UE) and a UPF) without being dropped.
@@
The algorithms for mapping the TSN configurations to 5G configuration values are current research topics.The work of [122] proposed mapping the default priority value of a 5G frame to the Priority Code Point (PCP). Other work by [123] consider that TSN frames shall be considered DC-GBR, while [121] propose the mapping of TSN traffic into 5G QoS 5QI characteristics. For example, the resource type characteristic is derived from the type of TSN flow: deadline constrained, jitter constrained or bandwidth constrained, per the truth table in table VII, with similar analysis for the other characteristics within 5QI.
TABLE VII: Mapping between TSN flow characteristics and 5QI values by [121].
@@
The integration of deterministic communications is not exclusive to TSN standards. DetNet standards can be used in conjunction with TSN standards to provide both upper and lower latency boundaries (while TSN can only ensure an upper bound), better scalability and interoperability with existing IP infrastructure (since it acts on OSI layer 3) [125]. Both sets of standards can act simultaneously, with [125] proposing that each TSN domain would be mapped into a subnet (or a tunnel) of a DetNet network, as depicted in fig. 11.
III-E3 In vehicular context
@@
Not only TSN has to be properly configured, but interoperability with other protocols must be assured. Ethernet with TSN must coexist and be interoperable between with other wired and wireless access technologies, including legacy automotive standards such as CAN. The synchronization strategies between the vehicular Ethernet backbone and CAN sync is current investigation by [19, 71], with three strategies to consider:
@@
IV Open challenges
Per the current state-of-the-art, it is essential to create tools to support a more agile service development, that considers the services requirements in terms of bandwidth, safety, criticality and timing of delivery of its information, while ensuring a interoperability and modularity to support (almost) any vehicle. These tools shall take into account not only the orchestration of those services throughout the vehicular zonal E/E architecture and its ECUs but also the network interconnecting those ECUs and connecting the vehicle to other elements on the smart city road. Therefore, determinism communications and its standards, tendentially and historically thought for static well-known environments, need to be able to be dynamically configured to support the highly dynamic environment within vehicles and their equally dynamic surroundings.
@@ IV Open challenges
Other challenges include:
@@
    Agility of service development: simplify and harmonize vehicular development processes — (almost) everything as software and software processes.
@@
    Data driven: vehicular architectures should be data driven in order to process great quantities of data from camera/lidar/sensors streams in use cases like autonomous driving [132].
@@
    Safety issues: following zero trust policies is essential.
@@
    Resource orchestration: leverage the unused resources — but not necessarily on a in-vehicle point of view, while considering costs and capabilities (CPU bound, I/O bound) from the hardware into the scheduling and orchestration.
@@
    Interoperability: multi-vendor, multi-hardware and multi vehicle interoperability.
@@
    Modularity: towards a more manageable stack where automakers don’t do everything.
@@
    Next-gen storage: consider cloud storage and solutions such as InterPlanetary File System (IPFS) towards use cases such as OTA updates.
@@
    Clustering: support the clustering of ECUs and the deployment of vehicular services across clusters, with reinforcement learning to choose the best node affinity.
@@
    Automatic generation of design architectures: computer generated architectures for zonal vehicular ECUs
@@
    Legacy compatibility: solutions to integrate ECUs interconnected with deterministic communications in a zonal SOA architecture.
@@
    Virtualization and automatic scheduling: reallocate workloads depending of resources with general purpose hardware and virtualized infrastructure.
@@
Considering the current state-of-the-art, the main challenges of TSN and deterministic networks within vehicular contexts include:
@@
    Integration of TSN within wireless networks: Some efforts are being done, but it is current work as discussed in section III-E2 [133].
@@
V Potential use cases
Vehicular services within an SDV context can benefit both intra-vehicular functions and occupants — both driver and passenger — as well as other vehicles and the overall smart city environment. Figure 14 lays out the main areas of use cases where the SDV paradigm may be useful.
Figure 14: Overview of the potential use cases.
With the SDV paradigm, the vehicle may send the information about its occupant to other subsystems, or vehicles through communications, which must be deterministic in order to ensure the information is obtained not too soon nor too late, therefore risking lives.
In all described use cases, the SDV paradigm supports some adaptability of the vehicle to the occupants’ characteristics and state to support some automated action (instead of a manual process). The occupants, either driving or passengers, can greatly benefit from services for customization of vehicular functions, protection, monitoring, and companionship.
When considering areas of use cases for SDV, one can consider both the drivers — for which most use cases, including autonomous driving, tend to focus on — and occupants. Both can benefit from vehicular technology that adapts to each occupant while increasing their autonomy and fostering their dignity. Software-Defined Vehicle can use the existing sensors and actuators on vehicles to support new services to support drivers and occupants in situations automakers could not predict during their design phase.
The following subsections present some use cases that are envisioned as possibilities to validate the system(s) to be researched — intra and inter-vehicle deterministic use cases and critical non-active safety use cases that require adaptability — both at software level but also at the network level (while ensuring reliability and bounded latency).
Table VIII provides an overview of the main requirements envisioned for the use cases. Use cases can be characterized by several types of requirements. The work of [134] define several vehicular use cases and their characteristics in terms of deadline (soft/hard and interval), data complexity (simple/complex and large) and pattern (discrete/continuous).
TABLE VIII: Main requirements for the possible use cases (based on [135], with adaptations)
@@ V Potential use cases
    considering in an urban situation
@@
[…]
All these use cases require $N$ time critical services and $K$ services, allowing to prove the need for orchestration and schedulability of intra- and inter-vehicular services. In all these use cases, a certain vehicular service can either be fully deployed on one or more node or divided through multiple nodes / cluster of vehicles, with both scenarios considering the cloud as an option.
Also common to all use cases is their need for time criticality. While the use cases do not necessarily focus on controlling the critical systems intra-vehicle such as power-train and braking subsystems, the are essential for other passive safety functionalities. For example, the systems to be researched are not going to actuate on the brakes for collision avoidance, but might want to warn cars behind to slow down because there is a disabled person in the car.
V-A Towards autonomous driving for all road situations

### v2

@@ 3.7.2 In wireless context
  • 2.
    Shadowing and attenuation: from other physical obstacles and phenomena (e.g. reflection, absorption, diffraction) [105].
  • 3.
    Interference: from other devices communicating on the same spectrum frequencies. This issue is more pronounced when unlicensed spectrum (e.g. 2.4 GHz) is used since its usage is uncontrolled and prone to additional noise from other uncontrollable entities.
  • 4.
@@
  • 6.
    Bounding an end-to-end latency: or its jitter by integrating scheduling on transmission by each Access Point (AP).
  • 7.
    Integration of reliability technique: while considering and taking maximum advantage of the main different channel characteristics.
  • 8.
@@
Achieving determinism within a wireless communication requires that all these challenges are properly addressed. While fading, interference, shadowing and attenuation are mostly unavoidable without changes to the physical environment, interference can be minimised. The usage of other parts of the spectrum, possibly licensed, with wider channels, and frequencies in multiple links working in conjunction (multi link operation) can further decrease the overall noise within the physical layer and therefore decrease the jitter of latencies, and increase the determinism of the communications [98]. In addition, a periodic analysis of the spectrum usage can support a decision on the best frequencies to be used.
Regarding the unpredictable time to access the wireless medium, existing listen-before-talk solutions such as Carrier-Sense Multiple Access (CSMA) consider that in the event of a collision, before re-transmission a random amount of time shall be waited. This creates an unpredictable, non-deterministic jitter on the transmission of information. Other channel access solutions must be considered in order to achieve higher determinism on the transmission of the frames through both a wired or wireless medium, such as dividing the channel per frequency, time or both (e.g. Frequency-Division Multiple Access (FDMA),  Time-Division Multiple Access (TDMA) and  Orthogonal FDMA (OFDMA), respectively) or adding scheduling algorithms to listen-before-talk solutions ( CSMA-Collision Avoidance (CSMA-CA)) [106]. In alternative, solutions that consider collisions to be solved at the receptor side (e.g.  Non-orthogonal Multiple Access (NOMA)) can also be considered.
Another promising solution is a coordination and scheduling of the access to the wireless medium to ensure not only the typical concerns of fairness or throughput but the fulfillment of a specific time of transmission, ensuring more determinism than the previous approaches [98]. Within wireless mediums, the usage of the IEEE 802.11be (Wi-Fi 7) concept of restricted Target Wake Time (rTWT) (originally introduced for power saving) can further decrease conflicts during transmission. By defining when a certain station cannot transmit — and therefore any transmission ends before the start of this time — it is possible to further decrease collisions between critical and less critical frames being transmitted simultaneously.
@@
Regarding the bounding of end-to-end latency and jitter, the integration of shaping techniques, namely TAS/802.11Qbv or ATS is possible. Solutions for integration include mapping the scheduling policies to each transmission queue within an AP and by adding to each frame fields with the flow identification and characteristics (delay bound, maximum packet size) to map the scheduling policies to the 802.11be MAC. The overall scheduling would require a full knowledge of the network topology.
@@
Finally, regarding the resource management, the need for some sort of centralization in decision making for the configuration of the network nodes — APs and switches — is supported both by TSN and DetNet standards. TSN IEEE 802.1Qcc standard define that 2 entities (the previously defined CUC and CNC) are responsible for monitoring and defining resource allocation and configurations [109], while the DetNet Reference Architecture [88] considers the existence of a Controller Plane responsible for “topology discovery and maintenance, packet route selection and instantiation or path fail-over mechanisms” [110]. An architecture should therefore consider the existence at least of a controller that configures and monitors the network nodes and the information flow.
All the previous points may benefit from additional performance performance with a deployment considering hardware acceleration, for example through the usage of Application-Specific Integrated Circuits or FPGAs [111]. FPGAs, by allowing the reconfiguration of the hardware to fit the needs of the application(s), would add to the hardware acceleration, the flexibility of deployment and reconfiguring of new wireless APs that combine the previous techniques accelerated by hardware. Such acceleration is essential to ensure the processing times are sufficiently small to allow determinism in the overall communication.
@@
  • 1.
    Timer: used by the scheduler and synchronized by an external Pulse Per Second (PPS) signal;
  • 2.
    Scheduler: operating at OSI level 2, decides which frames should be sent and when are they due;
  • 3.
    Custom MAC: contains information for controlling frames, with changes from the original 802.11 access scheme to avoid collision between real-time and non-real-time frames and supporting the necessary synchronization between clocks;
  • 4.
@@
While this work offers a promising solution, it lacks full testing in mobility scenarios, integration with TSN and integration with other wireless standards such as 5G-NR. However, such integration with cellular communication technologies is possible. 3GPP Releases 16-18 defends that, ensuring ultra reliable and low latency communications start at the radio level and that TSN standards should be added to the overall architecture, with parts of the 5G core acting as TSN bridges [109], one per 5G User Plane Function (UPF[113]. This allows a transparent integration of 5G systems with a TSN system but require a set of translators between the 5G control plane and TSN standards [114]. The configuration of the TSN network can then be based on the 5G configuration, with the mapping between the TSN network configurations [115](such as priorities or TAS gate opening times) and 5G QoS parameters, namely the 5G QoS Identifier (5QI), a scalar that represents characteristics such as [113, 116]:
  • 1.
    Resource Type: distinguishes between Guaranteed Bit Rate (GBR), Delay Critical GBR (DC-GBR) or non-GBR, where the first two types of flows can pre-allocate resources;
  • 2.
    Priority Level: indicates a priority in scheduling resources. Can be used to prioritize the allocation of resources to certain flows;
  • 3.
    Delay Budget: an upper bound of the time a packet can spend inside a 5G system (between the User Equipment (UE) and a UPF) without being dropped;
  • 4.
@@
Developing algorithms to map TSN configurations to 5G configuration values remains an active area of research. The work of [117] proposed mapping the default priority value of a 5G frame to the Priority Code Point (PCP). Other work by [118] consider that TSN frames shall be considered DC-GBR, while [116] propose the mapping of TSN traffic into 5G QoS 5QI characteristics. For example, the resource type characteristic is derived from the type of TSN flow: deadline constrained, jitter constrained or bandwidth constrained, per the truth table in table 7, with similar analysis for the other characteristics within 5QI.
Table 7: Mapping between TSN flow characteristics (on the left) and 5QI values (on the right)  [116].
@@
3.7.3 In vehicular context
@@
Not only TSN has to be properly configured, but interoperability with other protocols must be assured. Ethernet with TSN must coexist and be interoperable between with other wired and wireless access technologies, including legacy automotive standards such as CAN. Synchronization strategies between the vehicular Ethernet backbone and CAN are currently being investigated by [14, 67], with three approaches under consideration as depicted in fig. 13:
  • 1.
@@
The integration of deterministic communications in vehicles is not exclusive to TSN standards. DetNet standards can be used in conjunction with TSN standards to provide both upper and lower latency boundaries (while TSN can only ensure an upper bound), better scalability and interoperability with existing IP infrastructure (since it acts on OSI layer 3) [126]. Both sets of standards can act simultaneously, with [126] proposing that each TSN domain would be mapped into a subnet (or a tunnel) of a DetNet network, as depicted in fig. 14.
Integration with cloud infrastructure is also possible. For example, within [127] is proposed a full-fledged dynamic interference-aware cloud scheduling architecture for latency-sensitive workloads.
4 Open challenges
@@
Per the current state-of-the-art, it is essential to create tools to support a more agile service development, that considers the services requirements in terms of bandwidth, safety, criticality and timing of delivery of its information, while ensuring a interoperability and modularity to support (almost) any vehicle. These tools shall take into account not only the orchestration of those services throughout the vehicular zonal E/E architecture and its ECUs but also the network interconnecting those ECUs and connecting the vehicle to other elements on the smart city road. Therefore, determinism communications and its standards, tendentially and historically thought for static well-known environments, need to be able to be dynamically configured to support the highly dynamic environment within vehicles and their equally dynamic surroundings. Other challenges include supporting and improving:
Figure 16: Overview of the potential use cases.
@@
    Agility of service development: simplify and harmonize vehicular development processes — (almost) everything as software and software processes;
@@
    Data driven services: vehicular architectures should be data driven in order to process great quantities of data from camera/lidar/sensors streams in use cases like autonomous driving [128];
@@
    Safety concerns: following zero trust policies is essential, as well confidential computing concerns.
@@
    Optimized resource orchestration: leveraging the unused resources — but not necessarily on a in-vehicle point of view, while considering costs and capabilities (CPU bound, I/O bound) from the hardware into the scheduling and orchestration;
@@
    Interoperability: multi-vendor, multi-hardware and multi vehicle interoperability;
@@
    Modularity: towards a more manageable stack where automakers don’t do everything;
@@
    Next-gen distributed storage: considering cloud storage and solutions such as InterPlanetary File System (IPFS) towards use cases such as OTA updates;
@@
    Clustering: support the clustering of ECUs and the deployment of vehicular services across clusters, with reinforcement learning techniques (or equivalent) to choose the best node affinity;
@@
    Automatic generation of design architectures: computer generated architectures for zonal vehicular ECUs;
@@
    Legacy compatibility: solutions to integrate ECUs interconnected with deterministic communications in a zonal SOA architecture;
@@
    Virtualization and automatic scheduling: reallocate workloads depending of resources with general purpose hardware and virtualized infrastructure;
@@
Considering the current state-of-the-art, within the dynamic deployment of vehicular services focusing on the communication part with TSN and deterministic networks, the main challenges include:
@@
    Integration of TSN within wireless networks: Some efforts are being done, but it is current work as discussed in section 3.7.2 [129].
@@
5 Potential use cases
Figure 16 lays out the main areas of use cases where the SDV paradigm may be useful.
Even if vehicles will tend to become driver-less, autonomous vehicles may also take decades to be fully on the roads, and they will coexist with manual-driven cars. As such, autonomous vehicles shall consider the needs of the other occupants. It is therefore essential to think about vehicles, not only as means of transportation that may tend towards not requiring drivers, but also as adaptive objects, that suit themselves to the needs of occupants.
Examples of such needs would be the vehicle itself recognizing the needs or potential health issues, and providing personalized profiles and automatic adjustments on seating, cooling systems, driving profiles, infotainment screens, lighting, audio, and other vehicular subsystems that may support some sort of impairment. But this adaptation must be carried out flawlessly, since any failure can mean harm to someone, or even death.
Table 8: Main requirements for some possible use cases (based on [130], with adaptations).
@@ 5 Potential use cases
    considering an urban situation
@@
When considering areas of use cases for SDV, one can consider both the drivers — for which most use cases, including autonomous driving, tend to focus on — and occupants. Both can benefit from the vehicular technology that adapts to each occupant while increasing its autonomy. Software-Defined Vehicle can use the existing sensors and actuators on vehicles to support new services to support drivers and occupants in situations automakers could not predict during their design phase.
The following subsections present some use cases that are envisioned as possibilities to validate the system(s) to be researched — intra and inter-vehicle deterministic use cases and critical non-active safety use cases that require adaptability — both at software level and network level (while ensuring reliability and bounded latency). In all described use cases, the SDV paradigm supports some adaptability of the vehicle to the occupants’ characteristics and state to support some automated action (instead of a manual process). The occupants, either driving or passengers, can greatly benefit from services for customization of vehicular functions, protection, monitoring, and companionship.
5.1 Requirements and characterization of vehicular services
Table 8 provides an overview of the main requirements envisioned for the use cases. Use cases can be characterized by several types of requirements. The study in [132] defines several vehicular use cases and their characteristics in terms of deadline (soft/hard and interval), data complexity (simple/complex and large) and pattern (discrete/continuous).
During the overall definition and characterization of any vehicular service supporting an use case, it is essential to formulate the service and characterize its tasks in terms of those requirements, with particular emphasis on characteristics such as [19, 39, 133, 134]:
  • 1.
    Priority and criticality;
  • 2.
    Data specifications: such as format, size and real time guarantee requirements (periodicity or deadline for receiving/sending data);
  • 3.
    Expected computing location: edge vs cloud;
  • 4.
    Computing resources: either in terms of CPU, memory, storage, or graphical resources;
  • 5.
    Required real-time guarantee levels: in hard real-time, weakly hard real-time or soft real-time, depending on the impact of missed deadlines;
  • 6.
    Characteristics specific to the vehicular domain: such as the mobility of vehicular nodes and associated communication handovers and failures, and the heterogeneity and overall low specification of the ECUs;
  • 7.
    Non-functional requirements: noteworthy examples include high interoperability (including backwards compatibility with older vehicles and other vehicle brands), reliability, availability, scalability, extensibility, adaptability, testability and cost minimization.
These characteristics can be seen as the Service Level Agreement (SLA) of the service. Such SLAs should and could be defined by the developer of each vehicular service within a descriptor file for the service to support selecting the best technologies/stacks and configurations. Such configurations would be ensured/enforced by a distributed configurator through a set of QoS tools to enforce policies on the task scheduling and communication technologies tasks [134]. On task scheduling, the configurator would be responsible for a dynamic, intelligent, task scheduling and resource estimation, provisioning and allocation, including the offloading to other layers of communication (fog and cloud) while keeping the original non-functional requirements. Such scheduling could include the prediction of the overall resource usage while considering heterogeneity and computational cost of the nodes for optimization purposes [135, 136]. Estimation and prediction of the mobility of a certain service deployment based on trends of resource usage and prediction of the mobility of the vehicle would be essential in scenarios where the service requires deployment on multiple vehicles and/or external infrastructure to ensure coordination with other vehicles or entities [137].
Within the communication technologies, the configurator would be responsible for the dynamic management of the communication technologies during the mobility of the vehicle. Solutions would include the federation of multiple physical wireless and wired access layers in order to obtain a multi technology solution. Such management could also include slicing of the communication solution to support different levels of QoS specified by the service descriptors, as well as the management of the high mobility with the management of reactive and proactive handovers leveraging SDN [138] based solutions;
5.2 Services for autonomous driving in all road situations